# Local converse theorems and Langlands parameters

### v4

@@ 8. A variant
It is false in general that one can replace the normalizer of $r({}^{L}G)$ by $r({}^{L}G)$ itself. For example in Case (f) above, two admissible homomorphisms $\phi_{1}$ and $\phi_{2}$ with $\phi_{2}=s\phi_{1}s^{-1}$ for orthogonal $s$ with determinant $-1$ will satisfy the equality of twisted gamma factors, but might be inequivalent.
To conclude this paper, we make the following observation for classical groups when $F$ has characteristic zero. Assume the validity of the Gross-Prasad and Rallis conjecture and that the generic local Langlands correspondence $i$ of [JL3] coincides with that obtained by endoscopic transfer in [Art] and [Mok]. This latter fact might be known, and we refer to [AHKO, Theorem B.2 and Remark B.3] in the case of symplectic and split orthogonal groups. Then using [JL3, Theorem 6.20] and thanks to the uniqueness of generic representations inside L-packets with respect to a prescribed non degenerate character proved in [Ato], Theorem 8.2 coincides with the local converse theorems mentioned in the introduction, using Langlands-Shahidi gamma factors.

### v5

@@ 8. A variant
It is false in general that one can replace the normalizer of $r({}^{L}G)$ by $r({}^{L}G)$ itself. For example in Case (f) above, two admissible homomorphisms $\phi_{1}$ and $\phi_{2}$ with $\phi_{2}=s\phi_{1}s^{-1}$ for orthogonal $s$ with determinant 1-1- 1 will satisfy the equality of twisted gamma factors, but might be inequivalent.
To conclude this paper, we make the following observation for classical groups when $F$ has characteristic zero. Assume the validity of the Gross-Prasad and Rallis conjecture and that the generic local Langlands correspondence $i$ of [JL3] coincides with that obtained by endoscopic transfer in [Art] and [Mok]. This latter fact is true at least in the setting of generic supercuspidal representations of symplectic and split orthogonal groups, and we refer to [AHKO, Theorem B.2 and Remark B.3]. Then using [JL3, Theorem 6.20] and thanks to the uniqueness of generic representations inside L-packets with respect to a prescribed non degenerate character proved in [Var], Theorem 8.2 coincides with the local converse theorems mentioned in the introduction, using Langlands-Shahidi gamma factors.